OpenStudy (austinl):

A neutron consists of one “up” quark of charge +2e/3 and two “down” quarks each having charge -e/3. If we assume that the down quarks are 4.1 × 10^-15 m apart inside the neutron, what is the magnitude of the electrostatic force between them? Any thoughts?

OpenStudy (anonymous):

The question is wrong, but you can just use Coulomb's law to find the electrostatic force.

OpenStudy (austinl):

That is what I thought. Use q_1=+2e/3 q_2&q_3=-e/3 d=4.1E-15 m Any other thoughts beyond that?

OpenStudy (anonymous):

Oh, right, it's just asking about the force between the two negative charges. The positive one is irrelevant for this answer.

OpenStudy (austinl):

Okay, so only use q_2 and q_3 and the distance and plug those bad boys in and away I go?

OpenStudy (anonymous):

Yep

OpenStudy (austinl):

Okay, gimme just a minute and I will get back to you :)

OpenStudy (austinl):

I got 13.609 N, which when plugged in is evidently incorrect.

OpenStudy (anonymous):

I got 2 N

OpenStudy (austinl):

I just did it again and got the same thing. And for the record, 2 N is wrong as well.

OpenStudy (anonymous):

1.52 N, rather Rounding

OpenStudy (austinl):

That is correct, could you show me how you had it plugged in? I had: F=((8.99E9)(-(1.60E-19))^2)/(4.1E-15)^2

OpenStudy (anonymous):

divide by 3^2

OpenStudy (anonymous):

the charge is e/3 and you put in e

OpenStudy (austinl):

Oh, stupid mistake.... Thank you very much.